# Quadratic-form Optimal Transport

### v2

@@ 7 Conclusion
The new framework of quadratic-form optimal transport (QOT) is proposed, with the key feature that the transport cost is linear in $\pi\otimes\pi$. Due to the possible non-convex structure, the QOT problem is difficult to solve numerically. We prove fundamental properties of QOT and highlight cases with explicit solutions, summarized in Table 1. Compared to classic OT, QOT gives rise to two new and special optimal transport plans, the V-transport and the diamond transport. The latter is particularly interesting since it is not Monge, but serves as a universal minimizer of wide classes of QOT problems (Theorems 5.5 and 5.9), some of which are non-convex.
@@ Appendix A Quadratic programming formulation
to minimize $\mbox{to minimize\leavevmode\nobreak\ \leavevmode\nobreak\ }&\sum_{i,k\in[N],j,\ell\in[M]}c(x_{i},y_{j},x_{k},y_{\ell})\pi_{ij}\pi_{k\ell}\\ \mbox{over\leavevmode\nobreak\ \leavevmode\nobreak\ }&\pi\in\mathbb{R}_{+}^{N\times M}\\ \mbox{subject to\leavevmode\nobreak\ \leavevmode\nobreak\ }&\sum_{j\in[M]}\pi_{ij}=\mu_{i}\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ for all $i\in[N]$}\\ &\sum_{i\in[N]}\pi_{ij}=\nu_{j}\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ for all $j\in[M]$}.\end{aligned}$ (A.1)
@@
to minimize $\mbox{to minimize\leavevmode\nobreak\ \leavevmode\nobreak\ }&\bm{\pi}^{\top}C\bm{\pi}\\ \mbox{over\leavevmode\nobreak\ \leavevmode\nobreak\ }&\bm{\pi}\in\mathbb{R}_{+}^{NM}\\ \mbox{subject to\leavevmode\nobreak\ \leavevmode\nobreak\ }&\pi\mathbf{1}_{N}=\bm{\mu}\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }\pi^{\top}\mathbf{1}_{M}=\bm{\nu}.\end{aligned}$ (A.2)
@@ Appendix B Independent coupling is rarely a QOT minimizer
[…]
@@
Consider $\mu$ uniform on $\mathfrak{X}=\{0,1\}$ and $\nu$ uniform on $\mathfrak{Y}=\{0,1,2,3\}$, with cost function $c$ given by $c(1,0,1,0)=1$, $c(0,0,0,0)=1$, $c(1,1,1,1)=-1$, $c(0,1,0,1)=-1$, $c(1,2,1,2)=1$, $c(0,2,0,2)=1$, and zero otherwise. Denote the transition probabilities from $0\in\mathfrak{X}$ to $0,1,2,3\in\mathfrak{Y}$ respectively by $p,q,r,1-(p+q+r)$. It follows that the transition probabilities from $1\in\mathfrak{X}$ are 1/$21/2-p,1/2-q,1/2-r,p+q+r-1/$ 2. So, the total transport cost is
@@ C.1 Basic facts
In probabilistic terms, $\iint c_{\gamma}\,\mathrm{d}\pi\otimes\mathrm{d}\pi=\mathbb{E}\big{[}|Y-Y^{\prime}|e^{-\gamma|X-X^{\prime}|}\big{]}$ for $(X,Y),(X^{\prime},Y^{\prime})\buildrel\mathrm{law}\over{\sim}\pi$ iid. The intuition is that the cost function (C.1) measures the difference between $Y$ and $Y^{\prime}$ when $X$ and $X^{\prime}$ are close, and the parameter $\gamma$ controls how the distance between $X$ and $X^{\prime}$ is discounted. The QOT problem is then formulated as
to $\begin{split}\mbox{to minimize}\quad&\iint|y-y^{\prime}|e^{-\gamma|x-x^{\prime}|}\,\mathrm{d}\pi(x,y)\mathrm{d}\pi(x^{\prime},y^{\prime})\\ \mbox{subject to}\quad&\pi\in\Pi(\mu,\nu).\end{split}$ (C.2)
@@ C.2 The first limit case
to $\begin{split}\mbox{to minimize}\quad&\iint|y-y^{\prime}||x-x^{\prime}|\,\mathrm{d}\pi(x,y)\,\mathrm{d}\pi(x^{\prime},y^{\prime})\\ \mbox{subject to}\quad&\pi\in\Pi(\mu,\nu).\end{split}$
@@
[…]
@@ C.3 The second limit case: Weak OT and measures of association
As $\gamma\to\infty$, the double integral has a formal limit of $\mathbb{E}[|Y-Y^{\prime\prime}|]$, where $Y,Y^{\prime\prime}$ are conditionally iid on $X$ and $(X,Y)\buildrel\mathrm{law}\over{\sim}\pi$, which is closely connected to measures of association studied by Deb et al. [2020, 2024].
@@
[…]
@@
where $Y$ and $Y^{\prime}$ are iid and $Y$ and $Y^{\prime\prime}$ are conditionally iid given $X$. Under some additional assumptions on $(X,Y)$, Deb et al. [2020] showed that if $X$ and $Y$ are non-degenerate, then $\eta(X,Y)=0$ if and only if $X$ and $Y$ are independent, and $\eta(X,Y)=1$ if and only if $Y$ is a measurable function of $X$. Our next result, with a self-contained proof, implies the above conclusion on $\eta$. It assumes only the first moment condition on $Y$, much weaker than the conditions in Deb et al. [2020].
@@
For two random variables, we write the convex order relation $Z\leqslant_{\rm cx}W$ if $\mathbb{E}[h(Z)]\leqslant\mathbb{E}[h(W)]$ for all convex functions $h$ such that the two expectations are well-defined. Note that $g(U)\leqslant_{\rm cx}f(X,U)\buildrel\mathrm{law}\over{=}Y$. Moreover, $g(U)\buildrel\mathrm{law}\over{=}Y$ when $X$ and $Y$ are independent. By Furman et al. [2017, Theorem 4.5], the functional $X\mapsto\int_{0}^{1}(2t-1)Q_{\mu}(t)\,\mathrm{d}t$ is strictly increasing in convex order, where $X\buildrel\mathrm{law}\over{\sim}\mu$. Therefore, $\mathbb{E}[|Y-Y^{\prime}|]$ is maximized if and only if $g(U)\buildrel\mathrm{law}\over{=}Y$. Therefore, for the maximizer $(X,Y)$, $\mathbb{E}[f(X,U)\mid U]=g(U)=f(X,U)$ holds true, implying that $X$ and $f(X,U)$ are independent as $X$ and $U$ are independent. This shows that the independent coupling $\pi_{\mathrm{ind}}$ is the unique maximizer of $\mathbb{E}[|Y-Y^{\prime}|]$.
@@ C.4 Numerical approximations for the optimizers
The QOT problems with cost functions $c_{\gamma}$ and $-c_{\gamma}$ are not convex, indicating that exact solutions may be difficult to compute numerically, and hence we apply heuristic local search algorithms to solve for an optimal transport map. More precisely, we apply the metaheuristic improvement method with pair exchange neighborhood (see Section 3.2 of Çela [2013], or Section 8.2.3 of Burkard et al. [2012]) to a discretized version of the maximization problem with cost function (C.1), which is a quadratic assignment problem. The discretization procedure is justified by Proposition 3.5. The resulting matching may approximate the minimizers and maximizers of $\iint c_{\gamma}\,\mathrm{d}\pi\otimes\mathrm{d}\pi$.
[…]
@@
Figure 3: Optimal coupling with quadratic-form cost function $c_{\gamma}(x,y,x^{\prime},y^{\prime})=|y-y^{\prime}|e^{-\gamma|x-x^{\prime}|}$ for $\gamma\in\{0.3,2,6\}$, where $\mu$ and $\nu$ are the empirical measures of $50$ simulated points from $\mathrm{U}(0,1)$ and from $\mathrm{N}(0,1)$, respectively. The optimal coupling is reported in ranks. The support appears close to a $\lambda$shape. The numerical procedure is based on the heuristic improvement method with pair exchange neighborhood with $500$ iterations, initiated from the comonotone transport $\pi_{\mathrm{com}}$.
[…]
@@ Appendix D Extended discussions on unsolved questions
    Brenier’s theorem in classic OT states that if $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$, $\mu$ is absolutely continuous, and the cost is given by the squared Euclidean distance $\|x-y\|^{2}$, then the (unique) transport plan is Monge and induced by the gradient of a convex function (Brenier [1987]; see also Santambrogio [2015, Theorem 1.17] for a more general version). The analogous question in the QOT context remains very challenging. Recent studies on the (2,2)-GW cost (6.3) suggest the existence of optimal 2-maps (transport plans supported on the union of the graphs of two maps, such as $\pi_{\mathrm{dia}}$, $\pi^{\lambda}_{\mathrm{x}}$, and $\pi_{\rm v}$ with $x,y$ flipped) and the non-existence of Monge minimizers under certain assumptions including absolute continuity of $\mu$ (Dumont et al. [2024, Theorem 3.6]). Our closed-form results (Propositions 4.1 and 4.3, and Theorems 4.9, 5.5, and 5.9) provide evidence that the existence of optimal 2-maps might be a universal phenomenon for many QOT problems (but not all of them, in view of Figure 4). This phenomenon is also present in many other extensions or special cases of classic OT (mostly on the real line) such as the martingale optimal transport (Beiglböck and Juillet [2016, Corollary 1.6]), the directional optimal transport (Nutz and Wang [2022, Corollary 2.9]), and OT with concave costs (Gangbo and McCann [1996, Theorem 6.4]).
@@
      Figure 5: Plots of the quadratic-form optimal coupling $(X,Y)$ and its copula version, with cost function $e^{-((x-x^{\prime})^{2}+(y-y^{\prime})^{2})/2}$ and marginals $\mu$ uniformly distributed on $[-1,0]\cup[1/2,1]$ and $\nu$ uniform on $[-1,-1/2]\cup[0,1]$, both approximated by $80$ iid samples. Since the QOT problem is convex (Theorem 5.5), we apply the OSQP solver to find the optimal transport plan, illustrated with the heatmaps. The copula remains of diamond shape, but differs from the diamond copula which is perfectly symmetric.
@@ Appendix E Omitted proofs of results from Section 3
A standard argument using Prokhorov’s theorem shows that $\{\mu_{n}\}$ and $\{\nu_{n}\}$ are equi-tight, and hence $\{\pi_{n}\}$ is relatively compact; see the proof of Theorem 6.8 of Ambrosio et al. [2021]. Let $\pi\in\Pi(\mu,\nu)$ be a limit point of $\{\pi_{n}\}$. Theorem 2.8 of Billingsley [2013] then implies that $\pi_{n}\otimes\pi_{n}\to\pi\otimes\pi$ weakly. Since $c$ is continuous and satisfies (3.1), we have $\iint c\,\mathrm{d}\pi_{n}\otimes\mathrm{d}\pi_{n}\to\iint c\,\mathrm{d}\pi\otimes\mathrm{d}\pi$ (see Van der Vaart [2000, Theorem 2.20] and the example that follows). On the other hand, for any $\hat{\pi}\in\Pi(\mu,\nu)$, Sklar’s theorem (McNeil et al. [2015, Theorem 7.3]) implies that there exists a copula $C$ such that the cdf of $\pi$ is equal to $C(F_{\mu},F_{\nu})$, where $F_{\mu}$ is the cdf of $\mu$. Take $\pi_{n}^{\prime}$ specified by its cdf $C(F_{\mu_{n}},F_{\nu_{n}})$. We have $\pi_{n}^{\prime}\in\Pi(\mu_{n},\nu_{n})$ and $\pi_{n}^{\prime}\to\hat{\pi}$ weakly. Therefore,

### v4

@@ 7 Conclusion
The new framework of quadratic-form optimal transport (QOT) is proposed, with the key feature that the transport cost is linear in $\pi\otimes\pi$. Due to the possible non-convex structure, the QOT problem is difficult to solve numerically. We prove fundamental properties of QOT and highlight cases with explicit solutions, summarized in Table 1. Compared to classic OT, QOT gives rise to two new and special optimal transport plans, the V-transport and the diamond transport. The latter is particularly interesting since it is not Monge, but serves as a universal minimizer of wide classes of QOT problems (Theorems 6.5 and 6.9), some of which are non-convex.
@@
Acknowledgments
We thank Jose Blanchet, Job Boerma, Marcel Nutz, and Johannes Wiesel for helpful discussions, Shengtong Zhang for proposing a simple proof of Theorem 6.2 in the case of uniform marginals, and Facundo Mémoli for bringing up a series of relevant literature on the Gromov–Wasserstein distance. We also thank two referees for their helpful comments that improved this paper. R. Wang acknowledges financial support from the Natural Sciences and Engineering Research Council of Canada (RGPIN-2024-03728, CRC-2022-00141).
Declaration of interests
The authors have no competing interests to declare that are relevant to the content of this article.
@@ Appendix A Quadratic programming formulation
[…]
@@
$\begin{aligned} \mbox{to minimize\leavevmode\nobreak\ \leavevmode\nobreak\ }&\bm{\pi}^{\top}C\bm{\pi}\\ \mbox{over\leavevmode\nobreak\ \leavevmode\nobreak\ }&\bm{\pi}\in\mathbb{R}_{+}^{NM}\\ \mbox{subject to\leavevmode\nobreak\ \leavevmode\nobreak\ }&\pi\mathbf{1}_{N}=\bm{\mu}\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }\pi^{\top}\mathbf{1}_{M}=\bm{\nu}.\end{aligned}$ (A.2)
@@ Appendix B Independent coupling is rarely a QOT minimizer
    $\tilde{c}(x,y):=\frac{1}{2}\iint\left(c(x,y,x^{\prime},y^{\prime})+c(x^{\prime},y^{\prime},x,y)\right)\mathrm{d}\mu(x^{\prime})\,\mathrm{d}\nu(y^{\prime})=\varphi(x)+\psi(y),\quad\pi_{\mathrm{ind}}\text{-a.e.},$ (B.2)
@@
Consider $\mu$ uniform on $\mathfrak{X}=\{0,1\}$ and $\nu$ uniform on $\mathfrak{Y}=\{0,1,2,3\}$, with cost function $c$ given by $c(1,0,1,0)=1$, $c(0,0,0,0)=1$, c(1,1,1,1)=1c(1,1,1,1)=-1, c(0,1,0,1)=1c(0,1,0,1)=-1, $c(1,2,1,2)=1$, $c(0,2,0,2)=1$, and zero otherwise. Denote the transition probabilities from $0\in\mathfrak{X}$ to $0,1,2,3\in\mathfrak{Y}$ respectively by p,q,r,1(p+q+r)p,q,r,1-(p+q+r). It follows that the transition probabilities from $1\in\mathfrak{X}$ are 1/2p,1/2q,1/2r,p+q+r1/21/2-p,1/2-q,1/2-r,p+q+r-1/2. So, the total transport cost is
@@ C.1 Basic facts
In probabilistic terms, $\iint c_{\gamma}\,\mathrm{d}\pi\otimes\mathrm{d}\pi=\mathbb{E}\big{[}|Y-Y^{\prime}|e^{-\gamma|X-X^{\prime}|}\big{]}$ for $(X,Y),(X^{\prime},Y^{\prime})\buildrel\mathrm{law}\over{\sim}\pi$ iid. The intuition is that the cost function (C.1) measures the difference between $Y$ and $Y^{\prime}$ when XX and $X^{\prime}$ are close, and the parameter $\gamma$ controls how the distance between XX and $X^{\prime}$ is discounted. The QOT problem is then formulated as
$\begin{split}\mbox{to minimize}\quad&\iint|y-y^{\prime}|e^{-\gamma|x-x^{\prime}|}\,\mathrm{d}\pi(x,y)\mathrm{d}\pi(x^{\prime},y^{\prime})\\ \mbox{subject to}\quad&\pi\in\Pi(\mu,\nu).\end{split}$ (C.2)
@@ C.2 The first limit case
$\begin{split}\mbox{to minimize}\quad&\iint|y-y^{\prime}||x-x^{\prime}|\,\mathrm{d}\pi(x,y)\,\mathrm{d}\pi(x^{\prime},y^{\prime})\\ \mbox{subject to}\quad&\pi\in\Pi(\mu,\nu).\end{split}$
@@
Without loss of generality, $\delta\in(0,1)$. Observe the elementary inequality $|u+e^{-u}-1|\leqslant u^{1+\delta}$ for $u\geqslant 0$. [^7] To see this, observe that $u+e^{-u}-1\geqslant 0$. Let $f(u)=u^{1+\delta}+1-u-e^{-u}$. Clearly, $f(u)\geqslant 0$ for $u\geqslant 1$. For $u\in(0,1)$, $f^{\prime}(u)=(1+\delta)u^{\delta}+e^{-u}-1\geqslant(1+\delta)u^{\delta}-u\geqslant 0$. Therefore, $f(u)\geqslant 0$ for $u\in[0,1]$. It follows that uniformly in $\pi\in\Pi(\mu,\nu)$,
@@ C.3 The second limit case: Weak OT and measures of association
As $\gamma\to\infty$, the double integral has a formal limit of $\mathbb{E}[|Y-Y^{\prime\prime}|]$, where $Y,Y^{\prime\prime}$ are conditionally iid on XX and $(X,Y)\buildrel\mathrm{law}\over{\sim}\pi$, which is closely connected to measures of association studied by Deb et al. [2020, 2024].
@@
$\lim_{\gamma\to\infty}\int\frac{\gamma}{2}e^{-\gamma|u|}g(x+u,y)f(x+u)\,\mathrm{d}u=g(x,y)f(x)\quad\pi\text{-a.e.}$ (C.5)
@@
where $Y$ and $Y^{\prime}$ are iid and $Y$ and $Y^{\prime\prime}$ are conditionally iid given $X$. Under some additional assumptions on $(X,Y)$, Deb et al. [2020] showed that if XX and $Y$ are non-degenerate, then $\eta(X,Y)=0$ if and only if XX and $Y$ are independent, and $\eta(X,Y)=1$ if and only if $Y$ is a measurable function of $X$. Our next result, with a self-contained proof, implies the above conclusion on $\eta$. It assumes only the first moment condition on $Y$, much weaker than the conditions in Deb et al. [2020].
@@
For two random variables, we write the convex order relation $Z\leqslant_{\rm cx}W$ if $\mathbb{E}[h(Z)]\leqslant\mathbb{E}[h(W)]$ for all convex functions $h$ such that the two expectations are well-defined. Note that $g(U)\leqslant_{\rm cx}f(X,U)\buildrel\mathrm{law}\over{=}Y$. Moreover, $g(U)\buildrel\mathrm{law}\over{=}Y$ when XX and $Y$ are independent. By Furman et al. [2017, Theorem 4.5], the functional $X\mapsto\int_{0}^{1}(2t-1)Q_{\mu}(t)\,\mathrm{d}t$ is strictly increasing in convex order, where $X\buildrel\mathrm{law}\over{\sim}\mu$. Therefore, $\mathbb{E}[|Y-Y^{\prime}|]$ is maximized if and only if $g(U)\buildrel\mathrm{law}\over{=}Y$. Therefore, for the maximizer $(X,Y)$, $\mathbb{E}[f(X,U)\mid U]=g(U)=f(X,U)$ holds true, implying that XX and $f(X,U)$ are independent as XX and $U$ are independent. This shows that the independent coupling $\pi_{\mathrm{ind}}$ is the unique maximizer of $\mathbb{E}[|Y-Y^{\prime}|]$.
@@ C.4 Numerical approximations for the optimizers
The QOT problems with cost functions $c_{\gamma}$ and $c_{\gamma}$ are not convex, indicating that exact solutions may be difficult to compute numerically, and hence we apply heuristic local search algorithms to solve for an optimal transport map. More precisely, we apply the metaheuristic improvement method with pair exchange neighborhood (see Section 3.2 of Çela [2013], or Section 8.2.3 of Burkard et al. [2012]) to a discretized version of the maximization problem with cost function (C.1), which is a quadratic assignment problem. The discretization procedure is justified by Proposition 4.5. The resulting matching may approximate the minimizers and maximizers of $\iint c_{\gamma}\,\mathrm{d}\pi\otimes\mathrm{d}\pi$.
The minimizer of $\iint c_{\gamma}\,\mathrm{d}\pi\otimes\mathrm{d}\pi$ is known to be the comonotone coupling (Proposition C.1) when $\mu,\nu$ are in the same location-scale family, and hence we choose uniform and normal marginals, that is, $\mu=\mathrm{U}(0,1)$ and $\nu=\mathrm{N}(0,1)$. In Figure 3, we report the approximate minimizers (normalized by their ranks) with parameters $\gamma\in\{0.3,\,2,\,6\}$, obtained from the numerical scheme above. Each of them has an interesting “ $\lambda$ -shaped” support, clearly different from the comonotone coupling, or any other explicit coupling that we studied.
@@
Figure 3: Optimal coupling with quadratic-form cost function $c_{\gamma}(x,y,x^{\prime},y^{\prime})=|y-y^{\prime}|e^{-\gamma|x-x^{\prime}|}$ for $\gamma\in\{0.3,2,6\}$, where $\mu$ and $\nu$ are the empirical measures of $50$ simulated points from $\mathrm{U}(0,1)$ and from $\mathrm{N}(0,1)$, respectively. The optimal coupling is reported in ranks. The support appears close to a $\lambda$ -shape. The numerical procedure is based on the heuristic improvement method with pair exchange neighborhood with $500$ iterations, initiated from the comonotone transport $\pi_{\mathrm{com}}$.
For the maximizers of $\iint c_{\gamma}\,\mathrm{d}\pi\otimes\mathrm{d}\pi$, we do not know explicit forms even for the case $\mu=\nu=\mathrm{U}(0,1)$, so we consider these marginals in the numerical scheme. In Figure 4, we report the approximate maximizers with parameters $\gamma\in\{0.3,\,2,\,6\}$. Maximizers for smaller $\gamma$ appear closer to $\pi_{\mathrm{dia}}$ (thus reassuring Proposition C.2), and for larger $\gamma$ appear closer to $\pi_{\mathrm{ind}}$ (thus reassuring Proposition C.4). The support of the optimizer seems to be contained in a certain symmetric convex shape $E_{\gamma}$ in $0,1]^{2}$. As $\gamma$ increases, the support expands, and less mass is concentrated near the boundary of $E_{\gamma}$ but more mass in the interior of $E_{\gamma}$.
@@ Appendix D Extended discussions on unsolved questions
    Brenier’s theorem in classic OT states that if $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$, $\mu$ is absolutely continuous, and the cost is given by the squared Euclidean distance $\|x-y\|^{2}$, then the (unique) transport plan is Monge and induced by the gradient of a convex function (Brenier [1987]; see also Santambrogio [2015, Theorem 1.17] for a more general version). The analogous question in the QOT context remains very challenging. Recent studies on the (2,2)-GW cost (3.3) suggest the existence of optimal 2-maps (transport plans supported on the union of the graphs of two maps, such as $\pi_{\mathrm{dia}}$, $\pi^{\lambda}_{\mathrm{x}}$, and $\pi_{\rm v}$ with $x,y$ flipped) and the non-existence of Monge minimizers under certain assumptions including absolute continuity of $\mu$ (Dumont et al. [2024, Theorem 3.6]). Our closed-form results (Propositions 5.1 and 5.3, and Theorems 5.9, 6.5, and 6.9) provide evidence that the existence of optimal 2-maps might be a universal phenomenon for many QOT problems (but not all of them, in view of Figure 4). This phenomenon is also present in many other extensions or special cases of classic OT (mostly on the real line) such as the martingale optimal transport (Beiglböck and Juillet [2016, Corollary 1.6]), the directional optimal transport (Nutz and Wang [2022, Corollary 2.9]), and OT with concave costs (Gangbo and McCann [1996, Theorem 6.4]).
@@
      Figure 5: Plots of the quadratic-form optimal coupling $(X,Y)$ and its copula version, with cost function $e^{-((x-x^{\prime})^{2}+(y-y^{\prime})^{2})/2}$ and marginals $\mu$ uniformly distributed on $1,0]\cup[1/2,1]$ and $\nu$ uniform on $1,-1/2]\cup[0,1]$, both approximated by $80$ iid samples. Since the QOT problem is convex (Theorem 6.5), we apply the OSQP solver to find the optimal transport plan, illustrated with the heatmaps. The copula remains of diamond shape, but differs from the diamond copula which is perfectly symmetric.
@@ Appendix E Omitted proofs of results from Section 4
A standard argument using Prokhorov’s theorem shows that $\{\mu_{n}\}$ and $\{\nu_{n}\}$ are equi-tight, and hence $\{\pi_{n}\}$ is relatively compact; see the proof of Theorem 6.8 of Ambrosio et al. [2021]. Let $\pi\in\Pi(\mu,\nu)$ be a limit point of $\{\pi_{n}\}$. Theorem 2.8 of Billingsley [2013] then implies that $\pi_{n}\otimes\pi_{n}\to\pi\otimes\pi$ weakly. Since $c$ is continuous and satisfies (4.1), we have $\iint c\,\mathrm{d}\pi_{n}\otimes\mathrm{d}\pi_{n}\to\iint c\,\mathrm{d}\pi\otimes\mathrm{d}\pi$ (see Van der Vaart [2000, Theorem 2.20] and the example that follows). On the other hand, for any $\hat{\pi}\in\Pi(\mu,\nu)$, Sklar’s theorem (McNeil et al. [2015, Theorem 7.3]) implies that there exists a copula $C$ such that the cdf of $\hat{\pi}$ is equal to $C(F_{\mu},F_{\nu})$, where $F_{\mu}$ is the cdf of $\mu$. Take $\pi_{n}^{\prime}$ specified by its cdf $C(F_{\mu_{n}},F_{\nu_{n}})$. We have $\pi_{n}^{\prime}\in\Pi(\mu_{n},\nu_{n})$ and $\pi_{n}^{\prime}\to\hat{\pi}$ weakly. Therefore,